# Generation and Preservation of Large Entangled States on Physical Quantum Devices

### v1

@@ Abstract
As quantum technology advances and the size of quantum computers grow, it becomes increasingly important to understand the extent of quality in the devices. As large-scale entanglement is a quantum resource crucial for achieving quantum advantage, the challenge in its generation makes it a valuable benchmark for measuring the performance of universal quantum devices. In this work, we study entanglement in Greenberger-Horne-Zeilinger (GHZ) and graph states prepared on the range of IBM Quantum devices. We generate GHZ states and investigate their coherence times with respect to state size and dynamical decoupling techniques. A GHZ fidelity of is measured on a 32-qubit GHZ state, certifying its genuine multipartite entanglement (GME). We show a substantial improvement in GHZ decoherence rates for a 7-qubit GHZ state after implementing dynamical decoupling, and observe a linear trend in the decoherence rate of µs for up to qubits, confirming the absence of superdecoherence. Additionally, we prepare and characterise fully bipartite entangled native graph states on 22 IBM Quantum devices with qubit counts as high as 414 qubits, all active qubits of the 433-qubit Osprey device. Analysis of the decay of 2-qubit entanglement within the prepared states shows suppression of coherent noise signals with the implementation of dynamical decoupling techniques. Additionally, we observe that the entanglement in some qubit pairs oscillates over time, which is likely caused by residual ZZ-interactions. Characterising entanglement in native graph states, along with detecting entanglement oscillations, can be an effective approach to low-level device benchmarking that encapsulates 2-qubit error rates along with additional sources of noise, with possible applications to quantum circuit compilation. We develop several tools to automate the preparation and entanglement characterisation of GHZ and graph states. In particular, a method to characterise graph state bipartite entanglement using just 36 circuits, constant with respect to the number of qubits.
@@ I Introduction
Producing large-scale entangled states on many qubits represents an important frontier in the race to scale up physical quantum computers. In quantum computers, entanglement manifests as non-classical correlations between qubits, such that qubits involved in an entangled system cannot be described independently from each other [1, 2]. Entanglement has been found to play a significant role in achieving quantum advantage [3, 4, 5, 6], and many quantum information processing protocols explicitly rely on entanglement as a resource [7, 8, 9, 10]. Furthermore, entanglement over large arrays of qubits is essential in many fault-tolerant computation schemes [11, 12, 13]. Multipartite entanglement is known for its complex structure [1], and modern quantum devices have passed the number of qubits a classical computer can store an arbitrary quantum state on [14, 15]. Consequently, as noisy intermediate-scale quantum (NISQ) devices [16] continue to improve in both size and error rates, it becomes paramount to characterise the ability of a quantum computer to generate and maintain large entangled states in a scalable manner.
@@ II Generation and Decay of Greenberger-Horne-Zeilinger States
GHZ states and verifying genuine multipartite entanglement
@@
Prepare the -qubit GHZ state in the form as exemplified in Fig. 1b.
@@
(Optionally) Apply a refocusing -pulse (i.e. an -gate) on each qubit, producing the state .
@@
The GHZ coherence can then be calculated as where are the overlap signal amplitudes which can be obtained by applying a Fourier transform to :
@@
GHZ state embedding on physical topologies
@@
In this work, we develop a topology-agnostic GHZ state preparation scheme that constructs tree-type GHZ preparation circuits on heavy-hexagonal layouts with minimal depth. The method additionally involves selecting qubit subsets with optimized parameters such as low two-qubit gate error rates. The exact parameter or combination of parameters to optimise is specified by assigning corresponding weights to the edges of the graph. We initially considered only the CNOT error rate. Our algorithm is divided into two components. The first component embeds a GHZ circuit as a directed tree branching out from a single source qubit. We apply CNOT gates in parallel prioritizing least depth first and lowest two-qubit error rate second. The second component runs the first component algorithm multiple times, trialling each physical qubit or a subset of physical qubits as the source qubit. We then select the circuit with the least depth as the primary criterion and the lowest total cost parameter as the secondary criterion. This way, we can conveniently embed least-depth GHZ states on larger and larger quantum devices.
@@
Scalable quantum readout-error mitigation (M3)
@@
where is the noisy probability vector returned by the system, is the ideal probability vector in the absence of readout errors (but may include other errors), and is the calibration matrix, where corresponds to the probability of bit string being measured as bit string .
There are limitations to QREM, with the most significant one being the exponential scaling of classical resources required to solve for with respect to . There are a variety of approaches to overcome this challenge, which often involve approximating the calibration matrix by reducing it to a tensor product of single-qubit components [18, 42]. More recently, a measurement error mitigation method called M3 (Matrix-Free Measurement Error Mitigation) has demonstrated order-of-magnitude improvement in mitigation time over traditional methods [42]. M3 relies on two main optimizations: subspace reduction of the full calibration matrix based on noisy input bit strings, and implementation of a matrix-free, iterative method for solving the system of linear equations. The method involves directly computing matrix elements as
This error mitigation technique has some caveats. Firstly, M3 works natively with quasi-probability distributions which can contain negative elements. These non-physical probabilities arise from the finite sampling and, while still adding up to one, are incompatible with methods such as MQC. Therefore, a classical algorithm is used to efficiently convert the quasi-probabilities into the closest physical probability distribution under the -norm, and it runs in time [43]. Furthermore, the speedup of the mitigation process depends on the sparsity of the measured probability distributions. A sparser corresponds to a greater subspace reduction, which makes M3 optimal for GHZ states since they have only two measurement outcomes. However, scalable mitigation of readout errors comes at the expense of increased uncertainty for measurement outcomes. The mitigation overhead is given by:
@@
Verifying 32-qubit GHZ states on a 127-qubit Eagle Processor
@@ III Preserving GHZ States via Dynamical Decoupling
@C=0.8em @R=0.7em
[…]
@C=0.8em @R=0.5em
[…]
@C=0.8em @R=0.85em
[…]
@@
Hahn Echo and Periodic Dynamical Decoupling for 7-Qubit GHZ States
@@
Decay experiments are performed for 7-qubit GHZ states prepared on the ibmq_mumbai device consisting of 27 qubits and a quantum volume of 128. For free decays, we increment the circuit delay by 1 µs up to a maximum delay of 15 µs. For decays preserved with Hahn echo or PDD, we increment by 2 µs up to a maximum of 30 µs. For PDD, we implement control-pulses in 0.5 µs regular intervals. We obtain both unmitigated results and results with QREM-applied via the M3 protocol. The GHZ state populations, coherences and fidelities are measured across five independent experiments, where circuits are executed with 4196 shots each.
@@
Figure 7 displays the experimental results. As shown in Fig. 7a, neither Hahn echo nor PDD led to a marked improvement of GHZ population times. In fact, PDD appears to accelerate the decay of GHZ populations. The results are related to how the ground and excited state populations evolve with respect to relaxation errors on superconducting quantum devices. In detail, qubits in the excited state will eventually spontaneously decay into the ground state at a rate described by the relaxation time. After a sufficient amount of time, a quantum computer will reset to the all ground state (with some fluctuations). The decay in GHZ populations is primarily caused by bit flips in the state due to thermal relaxation, although environmental noise can also cause random bitflips. The application of control pulses in Hahn echo or PDD, which flip the ground and excited state probabilities, will do little to prevent relaxation errors. In fact, as shown in the PDD curve, repeated application of -gates only introduces additional noise from single-qubit gate errors.
@@
Figure 7b shows substantial improvement in GHZ coherence times in experiments with Hahn echo and PDD. The GHZ coherences quantify the non-classical correlations between the and states. For free decays, this correlation drops to effectively zero by around µs. In contrast, the PDD curve maintains a measured GHZ coherence of after µs. Hahn echo also appreciably prolongs GHZ coherence times, resulting in at µs, albeit to a lesser extent. These results highlight the efficacy of dynamical decoupling-based techniques in protecting GHZ states against dephasing errors (related to dephasing times) on superconducting quantum computers.
@@
The decay of GHZ state populations and coherences are combined in Fig. 7c, which plots the GHZ state fidelities as a function of circuit delay . Both Hahn echo and PDD are shown to be effective techniques for preserving GHZ fidelities. It is interesting to observe that although PDD is superior to Hahn echo for preserving GHZ coherences, it is worse at preserving GHZ populations. This is likely due to additional noise introduced from the PDD gate sequences, which has the effect of obfuscating population measurements in exchange for better protection against decoherence. As a result, Hahn echo exhibits roughly similar performance to PDD in preserving overall fidelities. In future experiments, it may be worth testing more advanced dynamical decoupling protocols such as UDD [49], which are shown to be more typically more effective than PDD [50]. We comment that the small differences in the initial values of , and are likely attributed to device drift. We also remark that applying QREM mainly increases the initial , and values with little influence to the decay rates. Next, we evaluate the scaling of GHZ decoherence rates as a function of state size.
Scaling of Decoherence Rates and GHZ Size
@@
Studying the decoherence rates of multipartite states as a function of their size may provide vital insight into the noise underlying a quantum system. The strength and nature of this noise can determine the feasibility of scaling up a quantum device. In particular, it may reveal whether a system exhibits superdecoherence. Superdecoherence describes the coupling of qubits to a single reservoir, which cause qubit decoherence rates to scale with the size of the system [27]. Such an effect is detrimental to the realisation of large-scale, fault-tolerant quantum computers. GHZ states are particularly convenient for detecting superdecoherence due to their high sensitivity to noise. In detail, GHZ states accumulate decoherence between qubits, so if the dominant noise model is uncorrelated across qubits (i.e. the decoherence rate per qubit is constant), we expect GHZ decoherence rates to scale linearly with the number of qubits. In contrast, if the dominant noise model is correlated across qubits (i.e. the system experiences superdecoherence), we expect GHZ decoherence rates to scale polynomially.
[…]
@@
GHZ Decoherence Scaling Results
Figure 9a plots the -qubit GHZ state coherences (normalized) as a function of circuit delay for on ibm_hanoi, which consists of 27 qubits and a quantum volume of 64. We increase total delay in increments of 2.5 µs. The maximum delay for each experiment ranges from µs for to µs for . For reference, the average CNOT gate time on ibm_hanoi is 385 ns. Data points represent the average measured coherence between five experiments, and error bars represent the standard deviation. We fit the plotted data with the exponential decay curve Eq. 10. We execute circuits with 4196 shots each. We mitigate readout errors using M3, although as shown in the previous section, this has little effect on the decay rates.
Most of the decay curves are modeled well by the exponential decay function. A notable exception, however, is the coherences, which appear to plateau slightly before exhibiting exponential decay. As a result, the decay fit incorrectly projects the initial coherence to be greater than 1. We observe a clear pattern of increasing decay rates with increasing GHZ size. In order to quantify this trend, we plot the decoherence rate and the GHZ size , shown in Fig. 9b. We take the error from the standard deviation, which we derive from the covariance matrix produced by the fitting algorithm. In agreement with Ozaeta & McMahon, we are able to well fit the data with the linear trendline µs with —now extending up to . (With an anomoalous data point at ). Our results support the notion that recent IBM Quantum transmon devices are naturally robust against superdecoherence. We remark that although we only measure coherence times for GHZ states of up to 15 qubits, MQC and dynamical decoupling-based techniques improve the initial values enough that one can feasibly extend the study to larger GHZ states, especially as error rates and coherence times improve.
@@
1 | - | - | |
2 | - | - | |
3 | |||
4 | - | - | |
5 | |||
6 | - | - | |
7 | |||
8 | - | - | |
9 | - | ||
11 | - | ||
13 | - | ||
15 | - |
We summarise the GHZ coherence times on ibm_hanoi in Table 1. For easy comparison, we include the measured GHZ coherence times from Ozaeta & McMahon’s experiments. For , we report an average increase of coherence times of 133% over previous results. Furthermore, we report longer coherence times on 15-qubit GHZ states than on previous 8-qubit GHZ states of μs and μs, respectively. The improvement in GHZ coherence times may be attributed to hard improvements in combination with superior coherence detection methods, incorporating dynamical-decoupling based techniques.
IV Efficient Bipartite Entanglement Characterisation in Whole-Device Graph States
Graph States
Graph states are a class of entangled multi-qubit states that are defined with respect to a connected graph. They are a generalisation of cluster states [51], and hence form a universal basis for measurement-based computation [52]. Graph states are additionally useful for quantum error correcting codes [53], quantum secure communication [54], quantum metrology [55], and probing Bell inequalities over multi-partite systems [56]. In quantum circuit notation, a graph state may be expressed as
where , is the set of edges connecting graph containing vertices (qubits), and represents a controlled-phase gate between adjacent qubits and . We can equivalently define the stabilising operator for each qubit in :

### v2

# Characterization of entanglement on superconducting quantum computers of up to 414 qubits

@@ Abstract
As quantum technology advances and the size of quantum computers grow, it becomes increasingly important to understand the extent of quality in the devices. As large-scale entanglement is a quantum resource crucial for achieving quantum advantage, the challenge in its generation makes it a valuable benchmark for measuring the performance of universal quantum devices. In this work, we study entanglement in Greenberger-Horne-Zeilinger (GHZ) and graph states prepared on the range of IBM Quantum devices. We generate GHZ states and investigate their coherence times with respect to state size and dynamical decoupling techniques. A GHZ fidelity of is measured on a 32-qubit GHZ state, certifying its genuine multipartite entanglement (GME). We show a substantial improvement in GHZ decoherence rates for a 7-qubit GHZ state after implementing dynamical decoupling, and observe a linear trend in the decoherence rate of µs-1 for up to qubits, confirming the absence of superdecoherence. Additionally, we prepare and characterize fully bipartite entangled native graph states on 22 superconducting quantum devices with qubit counts as high as 414 qubits, all active qubits of the 433-qubit IBM Osprey device. Analysis of the decay of 2-qubit entanglement within the prepared states shows suppression of coherent noise signals with the implementation of dynamical decoupling techniques. Additionally, we observe that the entanglement in some qubit pairs oscillates over time, which is likely caused by residual ZZ-interactions. Characterizing entanglement in native graph states, along with detecting entanglement oscillations, can be an effective approach to low-level device benchmarking that encapsulates 2-qubit error rates along with additional sources of noise, with possible applications to quantum circuit compilation. We develop several tools to automate the preparation and entanglement characterization of GHZ and graph states. In particular, a method to characterize graph state bipartite entanglement using just 36 circuits, constant with respect to the number of qubits.
@@ I Introduction
| Entangled State | Experiment | Device(s) | Section |
GHZ
GHZ state generation and verification up to 32 qubits |
ibm_washington
(127 qubits) |
Section II |
GHZ
Extending 7-qubit GHZ state lifetimes with dynamical decoupling |
ibmq_mumbai
(27 qubits) |
Section III.1 |
GHZ
Scaling of GHZ decoherence rate versus state size up to 15 qubits |
ibm_hanoi
(27 qubits) |
Section III.2 |
Graph state
Scalable whole-device bipartite entanglement characterization up to 414 qubits |
22 devices,
see Table 4 |
Section IV |
Graph state
Testing dynamical decoupling for 127-qubit whole-device bipartite entanglement |
ibm_brisbane
(127 qubits) |
Section V |
Producing large-scale entangled states on many qubits represents an important frontier in the race to scale up physical quantum computers. In quantum computers, entanglement manifests as non-classical correlations between qubits, such that qubits involved in an entangled system cannot be described independently from each other [1, 2]. Entanglement has been found to play a significant role in achieving quantum advantage [3, 4, 5, 6], and many quantum information processing protocols explicitly rely on entanglement as a resource [7, 8, 9, 10]. Furthermore, entanglement over large arrays of qubits is essential in many fault-tolerant computation schemes [11, 12, 13]. Multipartite entanglement is known for its complex structure [1], and modern quantum devices have passed the number of qubits a classical computer can store an arbitrary quantum state on [14, 15]. Consequently, as noisy intermediate-scale quantum (NISQ) devices [16] continue to improve in both size and error rates, it becomes paramount to characterize the ability of a quantum computer to generate and maintain large entangled states in a scalable manner.
@@
To enhance the paper’s navigability, we provide a guiding table in Table 1 summarizing all main experiments, the devices tested, and the relevant sections of the paper.
@@ II Generation and Decay of Greenberger-Horne-Zeilinger States
II.1 GHZ states and verifying genuine multipartite entanglement
@@
Prepare the -qubit GHZ state in the form as exemplified in Fig. 1(b).
@@
(Optionally) Apply a refocusing -pulse, i.e., an -gate on each qubit.
@@
The -qubit GHZ coherence can then be calculated as where are the overlap signal amplitudes which can be obtained by applying a Fourier transform to :
@@
II.2 GHZ state embedding on physical topologies
@@
In this work, we develop a topology-agnostic GHZ state preparation scheme that constructs tree-type GHZ preparation circuits on heavy-hexagonal layouts with minimal depth. The method additionally involves selecting qubit subsets with optimized parameters such as low two-qubit gate error rates. The exact parameter or combination of parameters to optimize is specified by assigning corresponding weights to the edges of the graph. We initially considered only the CNOT error rate. Our algorithm is divided into two components. The first component embeds a GHZ circuit as a directed tree branching out from a single source qubit. We apply CNOT gates in parallel prioritizing least depth first and lowest two-qubit error rate second. The second component runs the first component algorithm multiple times, trialling each physical qubit or a subset of physical qubits as the source qubit. We then select the circuit with the least depth as the primary criterion and the lowest total cost parameter as the secondary criterion. This way, we can conveniently embed least-depth GHZ states on larger and larger quantum devices.
@@
II.3 Scalable quantum readout-error mitigation (M3)
@@
where is the noisy probability vector returned by the system, is the ideal probability vector in the absence of readout errors (but may include other errors), and is the calibration matrix.
There are limitations to QREM, with the most significant one being the exponential scaling of classical resources required to solve for with respect to . There are a variety of approaches to overcome this challenge, which often involve approximating the calibration matrix by reducing it to a tensor product of single-qubit components [18, 42]. More recently, a measurement error mitigation method called M3 (Matrix-Free Measurement Error Mitigation) has demonstrated order-of-magnitude improvement in mitigation time over traditional methods [42]. M3 relies on two main optimizations: subspace reduction of the full calibration matrix based on noisy input bit strings, and implementation of a matrix-free, iterative method for solving the system of linear equations. For two bit-strings , where qubit 0 corresponds to the least significant bit, the method directly computes matrix elements as
where corresponds to the probability of the -th qubit being in state and measured in state . This error mitigation technique has some caveats. Firstly, M3 works natively with quasi-probability distributions which can contain negative elements. These non-physical probabilities arise from the finite sampling and, while still adding up to one, are incompatible with methods such as MQC. Therefore, a classical algorithm is used to efficiently convert the quasi-probabilities into the closest physical probability distribution under the -norm, and it runs in time [43]. Furthermore, the speedup of the mitigation process depends on the sparsity of the measured probability distributions. A sparser corresponds to a greater subspace reduction, which makes M3 optimal for GHZ states since they have only two measurement outcomes. However, scalable mitigation of readout errors comes at the expense of increased uncertainty for measurement outcomes. The mitigation overhead is given by:
@@
II.4 Verifying 32-qubit GHZ states on a 127-qubit Eagle Processor
@@ III Preserving GHZ States via Dynamical Decoupling
III.1 Hahn Echo and Periodic Dynamical Decoupling for GHZ States
@@
Decay experiments are performed for 7-qubit GHZ states prepared on the ibmq_mumbai device consisting of 27 qubits and a quantum volume [50] of 128. For free decays, we increment the circuit delay by 1 µs up to a maximum delay of 15 µs. For decays preserved with Hahn echo or PDD, we increment by 2 µs up to a maximum of 30 µs. For PDD, we implement control-pulses in 0.5 µs regular intervals. We obtain both unmitigated results and results with QREM-applied via the M3 protocol. The GHZ state populations, coherences and fidelities are measured across five independent experiments, where circuits are executed with 4196 shots each.
@@
Figure 7 displays the experimental results. As shown in Fig. 7(a), neither Hahn echo nor PDD led to a marked improvement of GHZ population times. In fact, PDD appears to accelerate the decay of GHZ populations. The results are related to how the ground and excited state populations evolve with respect to relaxation errors on superconducting quantum devices. In detail, qubits in the excited state will eventually spontaneously decay into the ground state at a rate described by the relaxation time. After a sufficient amount of time, a quantum computer will reset to the all ground state (with some fluctuations). The decay in GHZ populations is primarily caused by bit flips in the state due to thermal relaxation, although environmental noise can also cause random bitflips. The application of control pulses in Hahn echo or PDD, which flip the ground and excited state probabilities, will do little to prevent relaxation errors. In fact, as shown in the PDD curve, repeated application of -gates only introduces additional noise from single-qubit gate errors.
@@
Figure 7(b) shows substantial improvement in GHZ coherence times in experiments with Hahn echo and PDD. The GHZ coherences quantify the non-classical correlations between the and states. For free decays, this correlation drops to approximately 0.1 by µs. In contrast, the PDD curve maintains a measured GHZ coherence of after µs. Hahn echo also appreciably prolongs GHZ coherence times, resulting in at µs, albeit to a lesser extent. These results highlight the efficacy of dynamical decoupling-based techniques in protecting GHZ states against dephasing errors (related to dephasing times) on superconducting quantum computers.
@@
The decay of GHZ state populations and coherences are combined in Fig. 7(c), which plots the GHZ state fidelities as a function of circuit delay . Both Hahn echo and PDD are shown to be effective techniques for preserving GHZ fidelities. It is interesting to observe that although PDD is superior to Hahn echo for preserving GHZ coherences, it is worse at preserving GHZ populations. This is likely due to additional noise introduced from the PDD gate sequences, which has the effect of obfuscating population measurements in exchange for better protection against decoherence. As a result, Hahn echo exhibits roughly similar performance to PDD in preserving overall fidelities. In future experiments, it may be worth testing more advanced dynamical decoupling protocols such as UDD [49], which are shown to be more typically more effective than PDD [51]. We comment that the small differences in the initial values of , and are likely attributed to device drift. We also remark that applying QREM mainly increases the initial , and values with little influence to the decay rates. Next, we evaluate the scaling of GHZ decoherence rates as a function of state size.
III.2 Scaling of Decoherence Rates and GHZ Size
@@
Studying the decoherence rates of multipartite states as a function of their size may provide vital insight into the noise underlying a quantum system. The strength and nature of this noise can determine the feasibility of scaling up a quantum device. In particular, it may reveal whether a system exhibits superdecoherence. Superdecoherence describes the coupling of qubits to a single reservoir, which cause qubit decoherence rates to scale with the size of the system [27]. Such an effect is detrimental to the realization of large-scale, fault-tolerant quantum computers. GHZ states are particularly convenient for detecting superdecoherence due to their high sensitivity to noise. In detail, GHZ states accumulate decoherence between qubits, so if the dominant noise model is uncorrelated across qubits (i.e. the decoherence rate per qubit is constant), we expect GHZ decoherence rates to scale linearly with the number of qubits. In contrast, if the dominant noise model is correlated across qubits (i.e. the system experiences superdecoherence), we expect GHZ decoherence rates to scale polynomially.
@@
Figure 9(a) plots the -qubit GHZ state coherences (normalized) as a function of circuit delay for on ibm_hanoi, which consists of 27 qubits and a quantum volume of 64. We increase total delay in increments of 2.5 µs. The maximum delay for each experiment ranges from µs for to µs for . For reference, the average CNOT gate time on ibm_hanoi is 385 ns. Data points represent the average measured coherence between five experiments, and error bars represent the standard deviation. We fit the plotted data with the exponential decay curve Eq. 10. We execute circuits with 4196 shots each. We mitigate readout errors using M3, although as shown in the previous section, this has little effect on the decay rates.
Most of the decay curves are modeled well by the exponential decay function. A notable exception, however, is the coherences, which appear to plateau slightly before exhibiting exponential decay. As a result, the decay fit incorrectly projects the initial coherence to be greater than 1. We observe a clear pattern of increasing decay rates with increasing GHZ size. In order to quantify this trend, we plot the decoherence rate and the GHZ size , shown in Fig. 9(b). We take the error from the standard deviation, which we derive from the covariance matrix produced by the fitting algorithm. In agreement with Ozaeta & McMahon, we are able to well fit the data with the linear trendline µs-1 with —now extending up to . (With an anomoalous data point at ). Our results support the notion that recent IBM Quantum transmon devices are naturally robust against superdecoherence. We remark that although we only measure coherence times for GHZ states of up to 15 qubits, MQC and dynamical decoupling-based techniques improve the initial values enough that one can feasibly extend the study to larger GHZ states, especially as error rates and coherence times improve.
@@
| 1 | - | - | |
| 2 | - | - | |
| 3 | |||
| 4 | - | - | |
| 5 | |||
| 6 | - | - | |
| 7 | |||
| 8 | - | - | |
| 9 | - | ||
| 11 | - | ||
| 13 | - | ||
| 15 | - |
We summarize the GHZ coherence times on ibm_hanoi in Table 2. For easy comparison, we include the measured GHZ coherence times from Ozaeta & McMahon’s experiments. For , we report an average increase of coherence times of 133% over previous results. Furthermore, we report longer coherence times on 15-qubit GHZ states than on previous 8-qubit GHZ states of μs and μs, respectively. The improvement in GHZ coherence times may be attributed to hard improvements in combination with superior coherence detection methods, incorporating dynamical-decoupling based techniques.
IV Efficient Bipartite Entanglement Characterization in Whole-Device Graph States
IV.1 Graph States
Graph states are a class of entangled multi-qubit states that are defined with respect to a connected graph. They are a generalization of cluster states [52], and hence form a universal basis for measurement-based computation [53]. Graph states are additionally useful for quantum error correcting codes [54], quantum secure communication [55], quantum metrology [56], and probing Bell inequalities over multi-partite systems [57]. In quantum circuit notation, a graph state may be expressed as
where , is the set of edges connecting graph containing vertices (qubits), and represents a controlled-phase gate between adjacent qubits and . We can equivalently define the stabilizing operator for each qubit in :